rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2871 — FS#6832 — Cooling switch room 3

Attached to Project— Datacenters
Maintenance
Existing
CLOSED
100%
We are currently switching the original cooling system of
the room 3 to the new generation.

The changeover is performed in steps that will last several days.
Date:  Tuesday, 03 July 2012, 12:20PM
Reason for closing:  Done
Additional comments about closing:  The switches were done yesterday.
The room 3 is completely on the new cooling system.
Comment by OVH - Tuesday, 26 June 2012, 18:09PM

The working are progressing and will be finished this week.